OpenStudy (anonymous):

If x^3=(2x)^2 and x≠ 0, then x=?

OpenStudy (phi):

\[ x^3=(2x)^2 \] (2x)^2 means 2x * 2x or 2*2*x*x or 4x^2 so \[ x^3 = 4x^2 \] add -4x^2 to both sides \[ x^3 - 4x^2 =0 \] I would factor out x^2. what do you get ?

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

It would be 4, sorry